OpenStudy (anonymous):

When Sean left for college, his parents decided to move from their two thousand four hundred square-foot home into a smaller house. If the size of their house decreased by thirty percent, how many square feet is their new house?

OpenStudy (anonymous):

2400-100% x-70% x=2400*70/100=24*70=1680

OpenStudy (anonymous):

Well, if original was 2400, and 30% was taken off, that is saying 2400 x 0.3 = Y 2400 - Y = X (size of smaller house) Post the result on here